OpenStudy (anonymous):

Need help checking where I went wrong on this problem. 1 - (2/3) y = 9/5 - y/5 + 3/5 Problem 15 * (1 - (2/3) y) = (9/5 - y/5 + 3/5) * 15 Distributive Property 15 - 30 = 135 - 15y + 45 Added and Subtracted like terms -15 = 180 - 15y Subtracted 180 both sides 195 = -15y Divided by -15 y = -13 I got Y = -13 However WolframAlpha.com tells me the answer is Y = -3. Where did I go wrong in my work?

OpenStudy (nubeer):

check ur answer after u applied Distributive Property. u haven't multiplied correctly do it again and answer is -3 if u stil dont get let me know but try first

OpenStudy (nubeer):

@JamesR4494

OpenStudy (anonymous):

@nubeer I can't find my error :c

OpenStudy (nubeer):

ok lets try this way.. 15(1) - 15(2/3)y = 15 (9/5) -15(1/5)y 15(3/5).. can u see what i did? tell me what will u get after just multiplying

OpenStudy (anonymous):

15 - 30y = 135 - 15y 45 I see my error now, thanks. Let me see if I can finally solve this problem now.

OpenStudy (anonymous):

I can't seem to get the correct answer after I multiply. I tried 4 times and got the wrong answer all 4 times.

OpenStudy (anonymous):

Upto Distributive Property it is right.. There are mistakes in Added And Subtracted Like terms..

OpenStudy (anonymous):

On multiplying 15 to both sides you must get: \(15-10y = 27-3y + 9\) Or : \(15 - 27 - 9 = 10y - 3y\) \(7y = -21\) Divide by 7 both the side to get y..

OpenStudy (anonymous):

your problem is here, 15 - 30 = 135 - 15y + 45 Added and Subtracted like terms you forgot to divide the -30/3

OpenStudy (anonymous):

Thank you all for the help, I'm apparently so tired i'm making simple yet big mistakes.

OpenStudy (anonymous):

yes, thats a very true statement, I would certainly agree